OpenStudy (anonymous):

Written in slope-intercept form, the equation x + 2y = 4 is 2y = -x + 4 y = -1/2x + 4 y = -1/2x + 2

OpenStudy (anonymous):

slope intercept form is y = mx + b x + 2y = 4 (we want y term by itself, so subtract x from both sides) 2y = -x + 4 (now we need to divide by 2) y = (-x + 4)/2 or y = -1/2x + 2